OpenStudy (anonymous):

Which is the sum of the 2 smallest angles in a right triangle? A 30 degrees B 45 degrees C 90 degrees D 180 degrees E 360 degrees

OpenStudy (welshfella):

well the 2 small angles add up to 90 degrees (because the 3 angles of a trianggle add up to 180 degrees)

OpenStudy (anonymous):

so would 180 be the answer? That's what my guess was from the beginning I just needed a double check.

OpenStudy (welshfella):

As you may know a right angled triangle has one of the angles = 90 degrees.

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

ALL 3 add up to 180 degrees so that cant be the answer.

OpenStudy (anonymous):

so less

OpenStudy (anonymous):

it would be 45?

OpenStudy (welshfella):

yes In fact they add up to 180 - 90 = 90 degrees

OpenStudy (anonymous):

so 45 is my answer?

OpenStudy (welshfella):

no you need the SUM of the TWO smallest angles.

OpenStudy (anonymous):

This doesn't make sense to me. (I'm sorry I'm really bad at math)

OpenStudy (anonymous):

30 or 90 is really be only options.

OpenStudy (misssmartiez):

Chose all that applys or no?

OpenStudy (anonymous):

no just one

OpenStudy (misssmartiez):

90 + 30 + 30 (30 +30 = 90) = 180. A triangles measure needs to be added up to 180. So.

OpenStudy (misssmartiez):



OpenStudy (misssmartiez):

get it.

OpenStudy (anonymous):

So 30 because 30+30=90 and 90+90=180

OpenStudy (misssmartiez):

yes!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Could you help with one more?

OpenStudy (misssmartiez):

no problem. sure.

OpenStudy (anonymous):

Last year, Frank earned $7.48 per hour. This year the company decided to give him a 30% increase in his hourly rate. How much is Frank earning per hour this year?

OpenStudy (anonymous):

A $7.78 B $9.72 C $10.48 D $22.44 E $24.93

OpenStudy (misssmartiez):

0.30 X 7.48